OpenStudy (anonymous):

Five positive integers have a mean of 10, a median of 12 and only one mode, which is 12. What is the difference between the largest and smallest of these numbers?

OpenStudy (anonymous):

Start by saying that either three of the numbers equal 12, or two of them equal 12 (because the mode is 12).

OpenStudy (anonymous):

I was thinking of doing (a+b+c+d+e)/5= 10 _,12,10,12,_ =50

OpenStudy (ganpat):

 a + b + c + d = 60 - 12 a+ b + c + d = 48 smallest can be 0 .. and largest can be 38.. to get a mean of 10... so i guess 38 -0 = 38 ???